OpenStudy (anonymous):

Integral sqrt(4+3x) dx

OpenStudy (anonymous):

\[\int\limits_{}^{} \sqrt{4+3x} dx\]

OpenStudy (experimentx):

let 4 + 3x = u then 3dx = du or, \( 1/3 \int \sqrt{u} du\) = \( 1/3 u^{3/2}/(3/2) \) = \( 2/9 * (4+3x)^{3/2} \)

OpenStudy (anonymous):

can someone do it without u substitution if possible i struggle with that

OpenStudy (kropot72):

\[2(4+3x)^{3/2}\]

OpenStudy (kropot72):

I got my result by first applying the standard rule for integration with respect to (4+3x).Then I differentiated the result with respect to x and found the result was twice the required result. So my result should have been: \[1/2(4+3x)^{3/2}\]

OpenStudy (anonymous):

wouldnt the standard rule just lead you to (2/3)(4+3x)^(3/2)?

OpenStudy (kropot72):

Correct

OpenStudy (inkyvoyd):

@Redlinl , you might want to practice u substitution, because i'm pretty sure you'll see more of it.

OpenStudy (anonymous):

\[\int\limits \left( ax + b \right)^n dx = \frac{1}{a \left( n + 1 \right)}\left( ax + b \right)^{n + 1} + C\]\[\int\limits \sqrt{4 + 3x} dx = \int\limits \left( 3x + 4 \right)^{\frac{1}{2}} dx\]\[= \frac{1}{3\left( \frac{1}{2} + 1 \right)}\left( 3x + 4 \right)^{\frac{1}{2} + 1} + C\]\[= \frac{2}{9} \left( 3x + 4 \right)^{\frac{3}{2}} + C\]\[= \frac{2}{9}\left( 3x + 4 \right)\sqrt{3x + 4} + C\]

OpenStudy (anonymous):

@redlinl You're gonna need U-substitution for calculus later on. U-Sub is actually the simplest technique for integration compared to the stuff you learn later on.

OpenStudy (anonymous):

okay but exraven will i need to memorize that property or is there any way to prove that easily?

OpenStudy (inkyvoyd):

@Redlinl , that's the chain rule.

OpenStudy (inkyvoyd):

well, more like reverse chain rule of differentiation.

OpenStudy (anonymous):

you dont need to memorize it, u-substitution is the best way

OpenStudy (kropot72):

Then when you differentiate back you get:\[(4+3x)^{1/2}\times3=3\sqrt{(4+3x})\]

OpenStudy (anonymous):

could i just integrate it in the way i would to get the int by parts formula with the chain rule?

OpenStudy (inkyvoyd):

By the way, anyone wanna provide me with a site where I can learn u-substitution? I can't do it for life.

OpenStudy (inkyvoyd):

Red, just learn integration with u-sub

OpenStudy (inkyvoyd):

You aren't going to pass the course without it.

OpenStudy (kropot72):

Which is three times the original expression to be integrated.

OpenStudy (anonymous):

well im not in any course now im just preparing for it by learning from a book but i could also use a good site

OpenStudy (inkyvoyd):

Well, I can too, because I'm learning it from a book also, and it's an 80s textbook ;)

OpenStudy (inkyvoyd):

I think differentials are not explained very well.

OpenStudy (inkyvoyd):

The result is that I can't use them.

OpenStudy (anonymous):

yeah and what does du actually mean? d/dx u?

OpenStudy (kropot72):

So the correct result is as follows: \[2/9(4+3x)^{3/2}\] Sorry about my earlier errors:(

OpenStudy (anonymous):

You can manipulate differentials like algebraic terms. (du) just means an infintely small change in (u)

OpenStudy (anonymous):

so the du is the derivative of u?

OpenStudy (anonymous):

the way i think of it, du means delta u, and dx means delta x, so dx/du, means delta x/ delta u, which is just the slope or the derivative. Remember, when you're integrating, you're actually approximating the area of the curve by adding an infinite number of boxes and their area. The dx and du is the base of the box, multiplied by the f(x) which is their height, you get the area of the boxes.

OpenStudy (anonymous):

does the d actually stand for \[\Delta\] because that would make this a little easier

OpenStudy (anonymous):

Yea thats how i think of them. In the Reimann sum we had the capital Delta, and in the integral we had dx, so they basically both mean the base of the rectangles. If you ever get to Arc Length of a curve, they'll explain a little more about what dy/dx really means.

OpenStudy (anonymous):

\[\sqrt[3]{?3x-1=3}\]